4. The line y=0.15x+0.79

represents an estimate of the average cost of gasoline for each year. The line
0.11x-y=-0.85

estimates the price of gasoline in January of each year (Bureau of Labor Statistics, 2006).
a) Do you expect the lines to be intersecting, parallel, or perpendicular? Explain your reasoning.
b) Use the equations of the lines to determine if they are parallel. What did you find?
c) Did your answer to part b confirm your expectation in part a?

Found 2 solutions by palanisamy, Alan3354:
Answer by palanisamy(496)

The givens of the lines are,
y=0.15x+0.79 ...(1)
and 0.11x-y=-0.85
0.11x+0.85 = y
y=0.11x+0.85 ...(2)
The slope of the first line is = 0.15
The slope of the second line is = 0.11
Since the slopes are not equal, the lines are not parallel. So they will intersect.
Product of the slopes is (0.15)*(0.11) = 0.0165 (not equal to -1)
So the lines are not perpendicular.
(1)-(2)=> 0 = 0.04x-0.06
-0.04x=-0.06
x=(-0.06)/(-0.04)
x=3/2 =1.5
(1) => y = 0.15*1.5+0.79
y = 0.225+0.79
y = 1.015
The point of intersection is (1.5,1.015)


Answer by Alan3354(69443)

The line y=0.15x+0.79
represents an estimate of the average cost of gasoline for each year.
------------------
The line 0.11x-y=-0.85
estimates the price of gasoline in January of each year
--------------------
a) Do you expect the lines to be intersecting, parallel, or perpendicular? Explain your reasoning.
They will intersect, because the slopes are different.
----------------------
b) Use the equations of the lines to determine if they are parallel. What did you find?
They're not, the slopes are different.
-------------------
c) Did your answer to part b confirm your expectation in part a?
Yes